OpenStudy (anonymous):

Which equation is an identity? a)11 – (2v + 3) = –2v – 8 b)5w + 8 – w = 6w – 2(w – 4) c)7m – 2 = 8m + 4 – m d)8y + 9 = 8y – 3

jimthompson5910 (jim_thompson5910):

5w + 8 – w = 6w – 2(w – 4) 5w + 8 – w = 6w – 2w + 8 4w + 8 = 4w + 8 ... which is ALWAYS true so this is an identity I'll let you check the others